A satellite weighing 1600 kg revolves aroundt the Earth in an orbit radius three times that of the Earth.If the radius of the Earth is 6.4x10⁶m,what is the centripetal force acting on the satellite?


1
Expert's answer
2021-01-06T13:53:23-0500

The centripetal force acting on the satellite is actually a Newton's gravity force:


"F = G\dfrac{mM}{r^2}"

where "m = 1600kg" is the mass of the satellite, "M =6\times 10^{24}kg" is the mass of the Earth, "r = 3R" is the radius of the orbit, "R = 6.4\times10^6m" is the radius of the Earth, and "G = 6.7\times10^{-11}m^3/(kg\cdot c^2)" is the gravitational constant. Thus, obtain:


"F = G\dfrac{mM}{9R^2}\\
F = 6.7\times10^{-11}\cdot \dfrac{1600\cdot 6\times 10^{24}}{9\cdot (6.4\times 10^6)^2} \approx 1745N"

Answer. 1745 N.